OpenStudy (anonymous):

The product of two consecutive integers is 71 more than their sum. Find the integers.

OpenStudy (anonymous):

\[n(n+1)=n+n+1+71\] is a start

OpenStudy (lgbasallote):

(x)(x+1) = 71 + (x + x+ 1) x^2 + x = 71 + 2x + 1 x^2 - x -72 = 0 (x-9)(x+8) = 0 x = 9; x = -8 if x =9, x+1 = 10 if x = -8, x = -7